OpenStudy (agentnao):

Okay, if 3(2) is raised to i - 1 and i=1, then would 3(2) just be raised to the zero power? And then it wouldn't be a geometric series anymore! Am I doing this wrong?

OpenStudy (agentnao):

http://screencast.com/t/4wnxqMpq Here is the question.

OpenStudy (anonymous):

Why wouldn't it be a geometric series anymore? If you raise something to the zeroth power, it is equal to 1.

OpenStudy (agentnao):

Right, but the the number 3(2) wouldn't change anymore. And for a geometric sequence, it's supposed to increased or decrease by a common ratio.

OpenStudy (agentnao):

Or is i just the next number you start with and i changes every time up until 7?

OpenStudy (anonymous):

You multiply it by 2 each time.

OpenStudy (anonymous):

First term is 3, second is 6, etc.

OpenStudy (agentnao):

Wait. How did you get that?

OpenStudy (agentnao):

I thought the variable i was supposed to tell you which number of the sequence you start with and the number on top of the sigma tells you how far of the sequence you're supposed to add.

OpenStudy (anonymous):

Let i = 1, then the 1st term term is 3 * 2^0 = 3 * 1 = 3

OpenStudy (anonymous):

Let i = 1, then the 1st term term is 3 * 2^0 = 3 * 1 = 3

OpenStudy (anonymous):

Yes, youre supposed to sum i = 1 -> 7

OpenStudy (agentnao):

So then i will change for each number. It won't always be 1?

OpenStudy (anonymous):

Correct, i isn't affected by the expression you're summing, it just increases by 1 for each term until you reach your max number.

OpenStudy (agentnao):

Oh, okay. I get it! I guess I forgot that, but that makes sense. Thank you for being patient with me and being so helpful!

OpenStudy (anonymous):

You're welcome :)

OpenStudy (agentnao):

Oops! And one more question. It's the last one, I swear! Would the total end up being 55987?

OpenStudy (agentnao):

Well, that was definitely wrong. I did it out by hand because that seemed like too big of an answer and it totally was. Turns out it's actually 381! Thanks though :)

OpenStudy (anonymous):

Yep I got 381 too, but I did it in my head so I wouldn't take that as fact!

OpenStudy (agentnao):

Wow! You did that in your head?? Very impressive! That's quite a few numbers to be adding. At least they weren't that big though.

OpenStudy (anonymous):

There's actually 2 neat properties you can use so the only calcuations I needed to do was 3 * ( 2^7 - 1) which is quite doable in the head.